A toaster uses 1400 watts and takes 2.30 minutes to cook a piece of bread. If electricity costs
$0.16 / (kW•hr), then how many pieces of bread can be cooked for $0.05.
1
Expert's answer
2019-11-04T10:36:21-0500

Total energy used = "1.4\times\frac{2.3}{60}=0.0536\ kW\ h"

No. Of pieces of bread in 1 kW h= "\frac{1}{0.0536}=18.63\ pieces"

So, "0.16\$\" produces 18.63 pieces of bread

Then, "0.05\$" will produce "\frac{18.63\times0.05}{0.16}=5.82\ pieces"